# Thinking with Images for Multimodal Reasoning: Foundations, Methods, and Future Frontiers

### v2

@@ 6.2 Implementation Frameworks
Reinforcement learning offers a powerful mechanism for enabling LMMs to acquire and internalize “Thinking with Images” behaviors through interaction with environments. To support this capability at scale, several frameworks have been developed to support scalable and customizable RL workflows. VeRL (ByteDance Seed team and verl community, 2024) delivers scalable RLHF solutions by integrating FSDP and Megatron-LM backends to optimize throughput, and has gained widespread adoption in academic communities. To address high-performance distributed training requirements, OpenRLHF (Hu et al., 2024c) employs a Ray-based architecture augmented with vLLM acceleration and advanced GPU-sharing mechanisms, significantly improving training speed over alternative methods. For efficient scaling across large GPU clusters, ROLL (ROLL Team and Other ROLL Contributors, 2025) utilizes multi-role distributed architectures and integrates SGLang to support models with over 200 billion parameters via compositional sample-reward routing. In contrast to these optimization-focused frameworks, TRL (von Werra et al., 2020) integrates seamlessly with the Hugging Face ecosystem to facilitate rapid prototyping. To enhance research flexibility, RL4LMs (Ramamurthy et al., 2022) and TRLX (Havrilla et al., 2023) provide modular, customizable components specifically designed for alignment research and novel algorithm development. Despite the flexibility and scalability of these frameworks, they remain limited in supporting models to “Thinking with Images” through interactive learning. To address this limitation, recent approaches have extended general-purpose RL methods with specialized mechanisms tailored for visual reasoning. For example, OpenThinkIMG (Su et al., 2025a) introduces an end-to-end framework that learns tool orchestration policies, integrating disparate vision tools within a unified optimization loop. In contrast, VILASR (Wu et al., 2025a) proposes a novel “drawing” action space that enables direct image manipulation for spatial reasoning tasks. Building upon the VeRL, VisionReasoner (Liu et al., 2025b) implements a unified multi-task framework that learns to route queries to specialized perception heads, jointly optimizing both routing and execution policies within a single RL iteration.
@@ 6.3 Conclusion and Future Frontiers
  • Benchmarks for Visual Manipulation and Construction. Although current multimodal reasoning benchmarks are increasingly, only a few require multi-step reasoning and imagination about the image content to arrive at the final answer. This limits the ability to evaluate models capable of “Thinking with Images”. Therefore, future benchmarks could include tasks that require the manipulation and imagination of complex visual elements in the given images. For example, a model might be required to draw an auxiliary line on a geometry diagram or sketch a navigational path through a maze. Furthermore, evaluating a model’s capacity for complex, real-world multimodal reasoning is of great significance. Such evaluations would prompt models to generate a sequence of intermediate visual states, thereby advancing the development of world models.
@@ 7 Applications
LMMs capable of “Thinking with Images” transform a diverse range of applications, enabling a more nuanced understanding of visual information and more intuitive human-machine interaction. This section explores key domains where these advanced capabilities show significant promise.
@@ 7.1 Interactive Systems and User Interfaces
LMMs are increasingly applied to Graphical User Interfaces (GUIs), where agents must perceive on-screen content and plan actions in a human-like manner (Cheng et al., 2024b; Zheng et al., 2024d; Wang et al., 2024f; Qin et al., 2025; Xie et al., 2024c; Wu et al., 2024g; Gou et al., 2024). Recent work explores how models can “Thinking with Images” to understand interfaces and perform interaction tasks. Foundation models like CogAgent (Hong et al., 2024) utilize high-resolution vision encoders to recognize fine-grained UI elements directly from pixels. This approach achieves state-of-the-art results on GUI benchmarks and can outperform agents that rely on extracted HTML, demonstrating the power of direct visual perception. Similarly, MobileLMM (Liu et al., 2024b) employs a multi-stage curriculum to master mobile app navigation, maintaining an internal visual state to reason about on-screen affordances and determine subsequent actions.
A significant trend is the development of unified vision-language-action (VLA) models that operate end-to-end. These agents use only visual observations and closed-loop reasoning to interact with GUIs. For example, ShowUI (Lin et al., 2024) introduces a lightweight VLA model that processes screenshots and outputs UI actions through interleaved token modeling, achieving strong zero-shot accuracy. In a similar vein, UI-TARS (Qin et al., 2025) functions as a pure vision agent, executing actions based solely on screenshots, while its System-2 reasoning module handles task decomposition. Aguvis (Xu et al., 2024) further standardizes GUI interaction across platforms using only pixel inputs, training on a multimodal trajectory dataset and using self-generated reasoning steps to guide its decisions. These models exemplify closed-loop visual reasoning by observing the screen, interpreting its structure, and generating executable actions without textual UI representations.
A primary challenge for GUI agents is obtaining sufficient training data to cover diverse applications and user goals. To address this, several data acquisition strategies have been explored. Explorer (Pahuja et al., 2025) uses automated web exploration to generate a large dataset of task trajectories, significantly boosting web automation performance. TongUI (Zhang et al., 2025d) converts online GUI tutorials into multimodal training data via automatic extraction. PC Agent-E (He et al., 2025c) demonstrates that few-shot expert demonstrations, augmented with synthetic variations, can enable agents to achieve strong generalization and outperform large-scale baselines.
@@
The application of ”Thinking with Images” to GUIs marks a paradigm shift from reliance on structured data like HTML to direct, pixel-based reasoning. This enables agents to operate across diverse platforms such as web, desktop, and mobile environments where programmatic access is unavailable. The core advancement lies in creating a closed-loop cognitive process. Models observe the visual state of an interface, interpret its affordances, and generate the next action, often using an internal visual chain-of-thought to plan multi-step tasks. This iterative cycle of perception and action allows for real-time, multi-turn control. While the diversity of interfaces poses a significant data challenge, recent work demonstrates promising solutions through automated data generation from web exploration, conversion of online tutorials, and efficient learning from a few expert demonstrations. These developments are paving the way for more general and capable UI agents that interact with digital environments as humans do.
@@ 7.2 Embodied AI and Robotics
To operate effectively in unstructured physical environments, embodied agents such as robots must reason with visual inputs. Recent advances in LMMs are enabling visual thinking for action planning, from sensor-grounded language understanding to internal scene simulation (Driess et al., 2023; Fan et al., 2025b; Zhao et al., 2025; Zawalski et al., 2025; Zhou et al., 2025c). Pioneering work like PaLM-E (Driess et al., 2023) demonstrated that integrating language with visual and sensor inputs allows for positive knowledge transfer from web-scale pretraining to real-world robot manipulation. Subsequent models like Interleave-VLA (Fan et al., 2025b) empower robots to process interleaved visual-language prompts and generate executable action sequences in zero-shot settings, improving generalization to unseen objects. Similarly, VIMA (Jiang et al., 2023) advances general-purpose robot control by using multimodal prompts to support open-vocabulary manipulation, enabling robots to interact with novel objects through LMM-generated structured cues.
Beyond direct perception-to-action policies, many approaches incorporate a visual chain of thought to enhance planning. For example, CoT-VLA (Zhao et al., 2025) improves temporal reasoning by prompting the model to predict intermediate visual subgoals before generating an action, which achieves substantial performance gains in manipulation tasks. Embodied ECoT (Zawalski et al., 2025) introduces a multi-step reasoning framework that explicitly considers plans, object features, and robot state before action prediction, significantly improving policy generalization. Pushing this concept further, VPRL (Xu et al., 2025) conducts planning entirely through sequences of generated images, akin to mental sketches. In spatial navigation tasks, VPRL outperformed language-based methods, suggesting that image-only reasoning may be more effective for certain spatial problems.
@@ 7.3 AI for Science
LMMs support scientific inquiry by interpreting visual data from experiments, microscope slides, and satellite imagery (Yan et al., 2025; Zhu et al., 2025b; Tan et al., 2025; Burgess et al., 2025). However, a consistent challenge emerges across these domains: while general-purpose models can describe visual scenes, they fail to grasp the physical dynamics and domain-specific complexities crucial for scientific reasoning (Schulze Buschoff et al., 2025). This gap highlights the need for models that can “Thinking with Images” in a way that is grounded in scientific principles.
To address this, work has focused on domain-specific training and reasoning strategies. In physics, Balazadeh et al. (2024) trained an LMM on thousands of simulated block tower scenarios augmented with physics descriptors. This specialized training improved the model’s ability to predict physical outcomes, allowing it to outperform larger, general models. In scientific image analysis, similar challenges appear. Verma et al. (2024) found that LMMs struggled with fine-grained microscopy images, misidentifying structures in cluttered scenes experts could interpret. This suggests future models must move beyond surface-level pattern recognition. In remote sensing, GeoChat (Kuckreja et al., 2023) demonstrates a successful adaptation. By fine-tuning an LMM on satellite imagery, the model supports multi-turn dialogues for tasks like change detection and geospatial VQA, achieving strong zero-shot performance. These examples show that while current models lack human-level expertise, they offer a foundation for future multimodal scientific agents that can parse data, reason visually, and potentially generate simulations as part of their thought process (Bercovich et al., 2025).
@@
The application of “Thinking with Images” in science is in its nascent stages, yet it holds significant promise for automating data analysis and hypothesis generation. Current LMMs show foundational capabilities in interpreting diverse scientific imagery, including physics simulations, microscopy slides, and remote sensing data. However, a critical finding is that general-world knowledge is insufficient for specialized scientific domains. The performance of these models improves dramatically with domain-specific adaptation. Training on curated datasets, such as physics simulations or annotated satellite images, allows the models to learn the relevant scientific priors and visual patterns that general models miss. This underscores a key insight: for AI to become a valuable partner in science, it must not only see but also understand the visual world through the lens of specific scientific disciplines. Future progress will likely depend on developing models that can integrate this domain knowledge more deeply, enabling them to perform more complex visual reasoning and even generate novel visual hypotheses.
@@ 7.4 Healthcare and Medicine
In healthcare, LMMs are integrating visual data with language to support clinical decision-making across a wide range of tasks (Li et al., 2023b; Moor et al., 2023; Chen et al., 2024c; Xia et al., 2024c, d; Seyfioglu et al., 2024; Xia et al., 2025c; Zhu et al., 2025d; Nath et al., 2025; Lai et al., 2025; Lin et al., 2025; Dai et al., 2025a). Early systems demonstrated the potential for generating diagnostic reports from medical images. For instance, Wang et al. (2024g) proposed an interactive computer-aided diagnosis (CAD) system that combines vision models with a frozen LLM to produce detailed reasoning for chest X-ray analysis. This approach improved diagnostic accuracy and generated patient-friendly explanations. More generalist models like HealthGPT (Lin et al., 2025) unify diverse medical tasks such as radiology QA, report generation, and image synthesis within a single framework, demonstrating strong performance by jointly reasoning over visual and textual inputs.
A critical challenge in this high-stakes domain is ensuring model reliability and mitigating issues like hallucination. Several strategies have been developed to address this. To improve diagnostic accuracy, Guo and Terzopoulos (2025) used expert-informed prompts to guide the model’s reasoning, significantly reducing missed cases in chest X-ray diagnosis. For specialized fields like digital pathology, PathologyVLM (Dai et al., 2025b) employs a two-stage training strategy to learn visual hierarchies of cells and tissues, enabling zero-shot QA on gigapixel slides. Other approaches leverage structured knowledge; for example, Li et al. (2025c) developed an anatomical ontology-guided framework to direct multi-step reasoning. To mimic clinical workflows, Xia et al. (2025a) proposed MMedAgent-RL, a multi-agent framework that simulates collaboration between general practitioners and specialists to enhance accuracy. Further advancing evaluation standards, MedEBench introduces a collaborative framework that leverages the reasoning capabilities of LVLMs to assess text-guided image editing models on clinically grounded medical visual concepts (Liu et al., 2025e). These efforts collectively show a shift from static predictions to dynamic, explainable medical reasoning.
@@
In the medical domain, the “Thinking with Images” paradigm is evolving diagnostic AI from simple label prediction to complex clinical reasoning. LMMs are now capable of generating interpretable diagnostic reports by grounding textual analysis in visual evidence from medical scans. This represents a significant step towards more transparent and trustworthy AI. However, the high-stakes nature of healthcare magnifies the importance of reliability. Consequently, a key focus of current research is not just on capability, but also on control and safety. Promising strategies to enhance robustness include guiding the model’s reasoning with expert knowledge, specializing models for niche domains like pathology, and designing multi-agent systems that simulate clinical collaboration. The ultimate objective is to create AI systems capable of reasoning visually, ensuring that their outputs are not only accurate but also verifiable and consistent with established medical workflows, thereby functioning as dependable assistants for clinicians.
@@ 7.5 Education and Training
In education, the “Thinking with Images” paradigm enables new AI tutors that do more than provide textual answers; they can visually demonstrate concepts. By generating diagrams, highlighting key features, or sketching solutions in real-time, these systems function as interactive whiteboards, making abstract subjects like mathematics more intuitive. This dynamic visual dialogue adapts to a learner’s needs, offering a more engaging and personalized educational experience. The application extends to professional training, where models provide nuanced feedback by analyzing visual data from real-world scenarios like classroom videos. This capability for contextual visual analysis allows feedback previously only possible through human observation. These advances signal the emergence of visually-grounded educational agents blending image and language understanding to deliver more effective, interactive, and human-like learning support.

### v3

@@ 6.2 Implementation Frameworks
Reinforcement learning offers a powerful mechanism for enabling LMMs to acquire and internalize “Thinking with Images” behaviors through interaction with environments. To support this capability at scale, several frameworks have been developed to support scalable and customizable RL workflows on various LMMs. VeRL (ByteDance Seed team and verl community, 2024) achieves scalable RLHF solutions by integrating FSDP and Megatron-LM backends to optimize throughput, and has gained widespread adoption in academic and industry communities. To address high-performance distributed training requirements, OpenRLHF (Hu et al., 2024c) employs a ray-based architecture augmented with vLLM acceleration and advanced GPU-sharing mechanisms, significantly improving training speed over alternative methods. For efficient scaling across large GPU clusters, ROLL (ROLL Team and Other ROLL Contributors, 2025) utilizes multi-role distributed architectures with ray and integrates megatron-core and vLLM to acclerate model training and inference. In contrast to these training optimization frameworks, TRL (von Werra et al., 2020) integrates seamlessly with the Hugging Face ecosystem to facilitate rapid development. To enhance the flexibility of frameworks, RL4LMs (Ramamurthy et al., 2022) and TRLX (Havrilla et al., 2023) provide modular, customizable components specifically designed for alignment research and novel algorithm development. Despite the flexibility and scalability of these frameworks, they remain limited in supporting LMMs to “Thinking with Images” through interactive learning. To address this limitation, recent approaches have extended general-purpose RL methods with specialized mechanisms tailored for visual reasoning. For example, OpenThinkIMG (Su et al., 2025a) introduces an end-to-end framework that learns tool orchestration policies, integrating disparate vision tools within a unified optimization loop. In contrast, VILASR (Wu et al., 2025a) proposes a novel “drawing” action space that enables direct image manipulation for spatial reasoning tasks. Building upon the VeRL, VisionReasoner (Liu et al., 2025b) implements a unified multi-task framework that learns to route queries to specialized perception heads, jointly optimizing both routing and execution policies within a single RL iteration.
@@ 6.3 Conclusion and Future Frontiers
  • Benchmarks for Visual Manipulation and Construction. While the number of multimodal reasoning benchmarks is growing, few require the multi-step reasoning and imagination necessary to arrive at a final answer. This gap limits the ability to evaluate models truly capable of “Thinking with Images”. Therefore, future benchmarks could include tasks that require the manipulation and imagination of complex visual elements in the given images. For example, a model might be required to draw an auxiliary line on a geometry diagram or sketch a navigational path through a maze. Furthermore, evaluating a model’s capacity for complex, real-world multimodal reasoning is of great significance. Such evaluations would prompt models to generate a sequence of intermediate visual states, in guiding the development of world models.
@@ 7 Applications
LMMs capable of “Thinking with Images” transform a diverse range of applications, enabling a more in-depth understanding of visual information and more intuitive human-machine interaction. This section explores key domains where these advanced capabilities show significant promise.
@@ 7.1 Interactive Systems and User Interfaces
LMMs are increasingly applied to Graphical User Interfaces (GUIs), where agents must perceive on-screen content and plan actions in a human-like manner (Cheng et al., 2024b; Zheng et al., 2024c; Wang et al., 2024e; Qin et al., 2025; Xie et al., 2024c; Wu et al., 2024g; Gou et al., 2024). Recent work explores how models can “Thinking with Images” to understand interfaces and perform interaction tasks. Foundation models like CogAgent (Hong et al., 2024) utilize high-resolution vision encoders to recognize fine-grained UI elements directly from pixels. This approach achieves state-of-the-art results on GUI benchmarks and can outperform agents that rely on extracted HTML, demonstrating the power of direct visual perception. Similarly, MobileLMM (Liu et al., 2024b) employs a multi-stage curriculum to master mobile app navigation, maintaining an internal visual state to reason about on-screen affordances and determine subsequent actions.
A significant trend is the development of unified vision-language-action (VLA) models that operate end-to-end. These agents use only visual observations and closed-loop reasoning to interact with GUIs. For example, ShowUI (Lin et al., 2024) introduces a lightweight VLA model that processes screenshots and outputs UI actions through interleaved token modeling, achieving strong zero-shot accuracy. In a similar effort, UI-TARS (Qin et al., 2025) functions as a pure vision agent, executing actions based solely on screenshots, while its System-2 reasoning module handles task decomposition. Aguvis (Xu et al., 2024) further standardizes GUI interaction across platforms using only pixel inputs, training on a multimodal trajectory dataset and using self-generated reasoning steps to guide its decisions. These models exemplify closed-loop visual reasoning by observing the screen, interpreting its structure, and generating executable actions without textual UI representations.
A primary challenge for GUI agents is obtaining sufficient training data to cover diverse applications and user goals. To address this, several data acquisition strategies have been explored. Explorer (Pahuja et al., 2025) uses automated web exploration to generate a large dataset of task trajectories, significantly boosting web automation performance. TongUI (Zhang et al., 2025d) converts online GUI tutorials into multimodal training data via automatic extraction. PC Agent-E (He et al., 2025b) demonstrates that few-shot expert demonstrations, augmented with synthetic variations, can enable agents to achieve strong generalization and outperform large-scale baselines.
@@
The application of “Thinking with Images” to GUIs marks a paradigm shift from reliance on structured data like HTML to direct, pixel-based reasoning. This enables agents to operate across diverse platforms such as web, desktop, and mobile environments where programmatic access is unavailable. The core advancement lies in creating a closed-loop cognitive process. Models observe the visual state of an interface, interpret its affordances, and generate the next action, often using an internal visual chain-of-thought to plan multi-step tasks. This iterative cycle of perception and action allows for real-time, multi-turn control. While the diversity of interfaces poses a significant data challenge, recent work demonstrates promising solutions through automated data generation from web exploration, conversion of online tutorials, and efficient learning from a few expert demonstrations. These developments are paving the way for more general and capable UI agents that interact with digital environments as humans do.
@@ 7.2 Embodied AI and Robotics
To operate effectively in unstructured physical environments, embodied agents such as robots must reason with visual inputs. Recent advances in LMMs are enabling visual thinking for action planning, from sensor-grounded language understanding to internal scene simulation (Driess et al., 2023; Fan et al., 2025b; Zhao et al., 2025a; Zawalski et al., 2025; Zhou et al., 2025c). Pioneering work like PaLM-E (Driess et al., 2023) demonstrated that integrating language with visual and sensor inputs allows for positive knowledge transfer from web-scale pretraining to real-world robot manipulation. Subsequent models like Interleave-VLA (Fan et al., 2025b) empower robots to process interleaved visual-language prompts and generate executable action sequences in zero-shot settings, improving generalization to unseen objects. Similarly, VIMA (Jiang et al., 2023) advances general-purpose robot control by using multimodal prompts to support open-vocabulary manipulation, enabling robots to interact with novel objects through LMM-generated structured cues.
Beyond direct perception-to-action policies, many approaches incorporate a visual chain of thought to enhance planning. For example, CoT-VLA (Zhao et al., 2025a) improves temporal reasoning by prompting the model to predict intermediate visual subgoals before generating an action, which achieves substantial performance gains in manipulation tasks. Embodied ECoT (Zawalski et al., 2025) introduces a multi-step reasoning framework that explicitly considers plans, object features, and robot state before action prediction, significantly improving policy generalization. Pushing this concept further, VPRL (Xu et al., 2025) conducts planning entirely through sequences of generated images, akin to mental sketches. In spatial navigation tasks, VPRL outperformed language-based methods, suggesting that image-only reasoning may be more effective for certain spatial problems.
@@ 7.3 AI for Science
LMMs are beginning to support scientific inquiry by interpreting complex visual data from experiments, microscope slides, and satellite imagery (Yan et al., 2025; Zhu et al., 2025b; Tan et al., 2025; Burgess et al., 2025). However, a consistent challenge emerges across these domains. While general-purpose models can describe visual scenes, they often fail to grasp the underlying physical dynamics and domain-specific principles crucial for rigorous scientific reasoning (Schulze Buschoff et al., 2025). The “Thinking with Images” paradigm directly addresses this gap by empowering models with a deeper, more active form of visual cognition that is grounded in scientific principles.
This advanced capability allows models to move beyond simple description to active analysis. For example, in remote sensing, a model can think with images by invoking a specialized tool to precisely quantify deforestation, a functionality explored in systems like GeoChat which supports complex geospatial dialogues (Kuckreja et al., 2023). In fields like microscopy where experts can easily interpret cluttered scenes that challenge LMMs (Verma et al., 2024), a model thinking with images could generate code to segment and highlight specific cell types. This creates a new, clarified visual representation that serves as a verifiable intermediate step in its analytical process.
The most profound application lies in using visual imagination for simulation. In physics, for instance, a model can learn to simulate the outcomes of physical events, such as predicting the stability of a block tower (Balazadeh et al., 2024). This transforms the model from a passive observer into an active simulator that can conduct “virtual experiments”. While current models are still developing this expertise, they are laying the foundation for future scientific agents that do not just parse data, but actively reason and experiment with it visually to uncover new insights (Bercovich et al., 2025).
@@
The application of “Thinking with Images” in science is in its nascent stages, but it already holds significant promise for automating data analysis and hypothesis generation. Current LMMs show foundational capabilities in interpreting diverse scientific imagery, including physics simulations, microscopy slides, and remote sensing data. However, a critical finding is that general-world knowledge is insufficient for specialized scientific domains. The performance of these models improves dramatically with domain-specific adaptation. Training on curated datasets, such as physics simulations or annotated satellite images, allows the models to learn the relevant scientific priors and visual patterns that general models miss. This points out a key insight: for AI to become a valuable partner in science, it must not only see but also understand the visual world through the lens of specific scientific disciplines. Future progress will likely depend on developing models that can integrate this domain knowledge more deeply, enabling them to perform more complex visual reasoning and even generate novel visual hypotheses.
@@ 7.4 Healthcare and Medicine
In healthcare, LMMs are evolving from passive image interpreters to active partners in clinical decision-making (Li et al., 2023b; Moor et al., 2023; Chen et al., 2024c; Xia et al., 2024c, d; Seyfioglu et al., 2024; Xia et al., 2025b; Zhu et al., 2025d; Nath et al., 2025; Lai et al., 2025; Lin et al., 2025; Dai et al., 2025a). The “Thinking with Images” paradigm is central to this evolution, enabling models to generate dynamic and explainable multimodal medical reasoning rather than static predictions. Early systems demonstrated this by combining vision models with LLMs to produce detailed diagnostic reports for chest X-rays, improving both accuracy and patient communication (Wang et al., 2024f).
A key challenge in this high-stakes domain is ensuring reliability. “Thinking with Images” offers powerful new strategies to enhance trustworthiness. For instance, models can be guided by expert knowledge, such as using anatomical ontologies to direct a multi-step visual search across a medical scan (Li et al., 2025c; Guo and Terzopoulos, 2025). This structured approach mitigates hallucination and mimics clinical workflows. For highly specialized fields like digital pathology, models are learning to programmatically construct visual hierarchies of cells and tissues, enabling precise analysis of gigapixel slides (Dai et al., 2025b). These methods transform the model from a black box into a transparent reasoner. At the forefront of this domain, applications are beginning to model complex cognitive processes, either by simulating clinical consultations through multi-agent collaboration or by intrinsically generating visual hypotheses. Frameworks like MMedAgent-RL simulate a consultation between a generalist and specialist AI, using dialogue to refine a diagnosis (Xia et al., 2025a). This represents a collaborative form of visual thought. Furthermore, generalist models like HealthGPT are beginning to natively synthesize new visual content, such as generating a healthy organ scan for comparison, as part of a unified reasoning process (Lin et al., 2025). Efforts to evaluate these advanced capabilities are also emerging, with benchmarks like MedEBench assessing the fidelity of text-guided clinical image editing (Liu et al., 2025e).
@@
In the medical domain, “Thinking with Images” is transforming diagnostic AI from a predictive tool into a collaborative reasoning partner. This paradigm grounds abstract medical analysis in concrete visual evidence, enabling models to generate interpretable reports, follow anatomical search paths, and even simulate clinical consultations. The critical insight from this high-stakes field is that the primary frontier is not just capability, but verifiability and safety. Therefore, the most promising advancements involve constraining and guiding the model’s visual thought process with expert knowledge and structured workflows. The ultimate objective is to develop AI systems that reason visually with clinical rigor, functioning as dependable assistants whose diagnostic conclusions are not only accurate but also transparent and auditable.
@@ 7.5 Education and Training
In education, the “Thinking with Images” paradigm enables new AI tutors that do more than provide textual answers; they can visually demonstrate concepts. By generating diagrams, highlighting key features, or sketching solutions in real-time, these systems function as interactive whiteboards, making abstract subjects like mathematics more intuitive. This dynamic visual dialogue adapts to a learner’s needs, offering a more engaging and personalized educational experience. The application extends to professional training, where models provide detailed feedback by analyzing visual data from real-world scenarios like classroom videos. This capability for contextual visual analysis allows feedback previously only possible through human observation. These advances signal the emergence of visually-grounded educational agents blending image and language understanding to deliver more effective, interactive, and human-like learning support.